# Minimax-Regret Sample Selection in Randomized Experiments

### v1

@@ 4.3 Worst-Case Performance
Table 2 presented the minimax sample allocations as well as their worst-case performances under different decision paradigms. We first highlight that different decision paradigms result in significantly distinct sample selections. The Egalitarian approach allocates a substantial portion of the sample to the older group (i.e., the group with a larger anticipated variance and a smaller weight). In contrast, both the minimax and the proportional selections prioritize the younger group that possesses a large weight, with the latter assigning it a higher number of samples. Moreover, it is evident that the minimax-regret sample selections are indeed regret minimax, and they consistently outperform the other sample allocations with respect to the specific worst-case scenarios defined by their targeting decision paradigms. Meanwhile, extreme selections that allocate all samples to one group result in unbounded worst case regret. Furthermore, when the policy-maker expects to make a single decision for the entire population, any sample selection other than proportional selection may result in arbitrarily bad decisions. Nevertheless, proportional selection’s performance deteriorates when group-level decisions are permitted, particularly when the policy-maker prioritizes the utility in the worst-off group.
@@ 4.4 Performance under Reported Incidence Rates
In addition to the worst-case performances, we also evaluate the performance of the minimax sample selections under the incidence rate reported in Baden et al. (2021). Again, we consider a data-generating distribution as stated in Theorem 2.1 where $Y_{g,i}(0)$ and $Y_{g,i}(1)$ are generated i.i.d. from Gaussian distributions $N(b_{g}-\tau_{g}/2,s_{0,g}^{2})$ and $N(b_{g}+\tau_{g}/2,s_{1,g}^{2})$, respectively. We obtain $\tau_{g}$, $s_{0,g}$ and $s_{1,g}$ through the incidence rates in Baden et al. (2021). Table 3 gives a summary of the incidence rates of severe COVID-19 and severe solicited adverse reactions reported in Baden et al. (2021). Assuming that the potential outcomes are independent, a normal approximation yields the parameters given in Table 4.
@@
Minimax (6100, 3218) 0.67 0.005 0.78 (6102, 3218) 1.16 1.73 2.26
Proportional (7734, 1584) 0.75 0.047 2.36 (7736, 1584) 1.07 0.68 3.03
Egalitarian (2068, 7250) 1.83 0.003 2.19 (2074, 7246) 2.16 6.07 2.34
@@
Table 5 summarizes the sample allocations and their performances. Despite the approximated noise levels used for guiding the selections being largely misspecified, the selections we investigated still achieve decent performances. In particular, although the minimax selection may appear conservative, it tends to outperform the proportional selection in most cases in this non-adversarial real-data example. In Case 1 where the minority group exhibits a higher signal-to-noise ratio, the minimax selection that allocates more equally outperforms under separate decisions; When only a joint decision is allowed, the Egalitarian selection that allocates a large portion of the sample to the easier-to-learn group achieves a smaller regret. Nevertheless, in Case 2 where the optimal decisions for the groups differ, the Egalitarian selection that aims to minimize regret in the worst-off group performs considerably worse, and it appears to be less robust to model misspecification in the setting we consider.
@@ 5 Discussion
While utilitarian social welfare is a natural and common choice, it may inadvertently favor groups with large proportion weights, potentially leading to unfair choices. An alternative is to consider the Egalitarian rule that suggests making decisions that maximize the minimum utility of all individuals in society under the least undesirable condition. Our findings demonstrate that this indeed results in equal sample allocation across all groups when the signal-to-noise ratio is the same across all groups. In scenarios where signal-to-noise ratios vary among groups, the resulting allocation suggests sampling proportionally to the variance of the treatment-control difference, ensuring an equal level of confidence in estimating the treatment effect across all groups.
@@
In this paper, we considered welfare-optimizing sample selection through the lens of minimax-regret decision rules. This approach is attractive in terms of its robustness and generality. However, in settings where prior evidence about the effectiveness of a treatment is available, it may be desirable to instead use Bayesian methods that incorporate this prior knowledge. Another interesting direction would be to investigate optimal sample selection under non-linear objectives. Models for cooperative bargaining can be used to motivate targeting expected log-utility (Nash, 1950; Kaneko and Nakamura, 1979); and Baek and Farias (2021) recently studied implications of targeting such objectives in bandit experiments. Finally, it may also be of interest to consider a richer welfare model where risks and benefits for study participants (who risk direct harms from trying out a potentially dangerous treatment) and the general population (whose risks and benefits are mediated by the overall study findings). Such extensions are beyond the scope of the present paper; however, we intend to study them in follow-up work.
@@ Acknowledgment
Appendix A Proof of Theorems
A.1 Proof of Theorem 16

### v2

@@ 4.3 Worst-Case Performance
Table 2 presented the minimax sample allocations as well as their worst-case performances under different decision paradigms. We first highlight that different decision paradigms result in significantly distinct sample selections. The egalitarian and the Neyman approaches allocate a substantial portion of the sample to the older group (i.e., the group with a larger anticipated variance and a smaller weight), with the egalitarian selection being more extreme. In contrast, both the minimax and the proportional selections prioritize the younger group that possesses a large weight, with the latter assigning it a higher number of samples. Moreover, it is evident that the minimax-regret sample selections are indeed regret minimax, and they consistently outperform the other sample allocations with respect to the specific worst-case scenarios defined by their targeting decision paradigms. Meanwhile, extreme selections that allocate all samples to one group result in unbounded worst case regret. Furthermore, when the policy-maker expects to make a single decision for the entire population, any sample selection other than proportional selection may result in arbitrarily bad decisions. Nevertheless, proportional selection’s performance deteriorates when group-level decisions are permitted, particularly when the policy-maker prioritizes the utility in the worst-off group.
@@ 4.4 Performance under Reported Incidence Rates
In addition to the worst-case performances, we also evaluate the performance of the minimax sample selections under the incidence rate reported in Baden et al. (2021), reproduced in Table 3. We then use Gaussian approximation to obtain the distribution of the difference-in-means treatment effect estimator, resulting in parameters reported in Table 4.
@@
Minimax (6100, 3218) 0.67 0.005 0.78 (6102, 3216) 1.16 1.73 2.26
Proportional (7734, 1584) 0.75 0.047 2.36 (7734, 1584) 1.07 0.68 3.03
Egalitarian (2068, 7250) 1.83 0.003 2.19 (2074, 7244) 2.16 6.07 2.34
Neyman (3244, 6074) 1.26 0.002 1.50 (3248, 6070) 1.70 5.19 1.75
@@
Table 5 summarizes the sample allocations and their performances. Despite the approximated noise levels used for guiding the selections being largely misspecified, the selections we investigated still achieve decent performances. In particular, although the minimax selection may appear conservative, it tends to outperform the proportional selection in most cases in this non-adversarial real-data example. In Case 1 where the minority group exhibits a higher signal-to-noise ratio, the minimax selection that allocates more equally outperforms under separate decisions; when only a joint decision is allowed, the egalitarian and the Neyman selections that allocate a large portion of the sample to the easier-to-learn group achieve a smaller regret. Nevertheless, in Case 2 where the optimal decisions for the groups differ, the egalitarian selection that aims to minimize regret in the worst-off group performs considerably worse, and it appears to be less robust to model misspecification than the Neyman selection in the setting we consider.
5 Bayes-Optimal Sample Selection
In Section 2, we identified the minimax-regret sample selection by minimizing the expected regret with $\tau$ chosen adversarially. The minimax rule can also be interpreted as a Bayes-optimal rules with a least favorable prior on $\tau$ (Wasserman, 2013). While the minimax rule has attractive worst-case guarantees, a decision-maker with meaningful information about $\tau$ may prefer to conduct Bayes-optimal sample selection using their chosen prior rather than with the least favorable one underlying the minimax rule. In this section, we briefly consider Bayesian sample selection under an informative prior, and examine alignment of the minimax-regret sample selections with Bayes-optimal ones in a simple example.
Within a Bayesian framework, the parameters are regarded as random variables governed by a prior distribution. Unlike in non-Bayesian approaches where the treatment effect $\tau$ is treated as an unknown fixed value, the experimenter holds a belief captured by a prior distribution $p$ that reflects their understanding of $\tau$ based on various sources such as historical data, literature, or common beliefs. [^6] In the Bayesian framework, we assume knowledge of all other components of the data-generating distribution. However, it is also possible to extend our framework to incorporate priors on these components. Below, we give a formal definition of the Bayes-optimal sample selection.
Definition 4.
Given a prior $p$ on the unknown parameter $\tau$, let $\delta^{p}(D)$ be a decision that is made according to the sign of the posterior mean of $\tau$, i.e.,
$\begin{split}&\delta^{p}(D)=I(\hat{\tau}^{\text{PM}}\geq 0),\\ &\hat{\tau}^{\text{PM}}=\mathbb{E}\left[\tau\,\big{|}\,D\right].\end{split}$ (26)
The Bayesian expected regret of a sample selection rule $\mathbf{n}\in\mathcal{N}$ subject to this prior is
$H^{p}(\mathbf{n})=\mathbb{E}_{\tau\sim p}\left[\mathbb{E}_{D}\left[R(\mathbf{n},\delta^{p})\,\big{|}\,\tau\right]\right].$ (27)
Such a sample selection is Bayes-optimal with respect to prior $p$ if $\mathbf{n}\in\operatorname{argmin}_{\mathbf{n}\in\mathcal{N}}H^{p}(\mathbf{n})$.
An advantage of the Bayesian decision-theoretic framework is its flexibility in modeling relationships between different groups. For example, one can account for information sharing across groups by adjusting the prior correlation between $\tau_{g}$’s. A high correlation implies that the treatment effects across groups are likely similar, and thus insights gained from one group can inform the others.
To obtain the Bayesian expected regret, both a prior and a data-generating distribution must be specified. When the prior is conjugate to the data-generating process, $\hat{\tau}^{\text{PM}}$ can be derived in closed form. Otherwise, computational methods such as Markov Chain Monte Carlo (Gelman et al., 2013) are needed to approximate $\hat{\tau}^{\text{PM}}$. In the supplementary material, we provide derivations of the posterior distribution to facilitate computation when the prior $p$ and the data-generating distribution $\mathcal{D}$ form a normal-normal conjugate.
To get some insight about the effect of prior choice on sample selection, we consider the following synthetic two-group setting. For unit $i$ from group $g$, $g=1,2$, we consider the following data-generating distribution:
$\begin{split}&Y_{g,i}(0)\sim N(b_{g}-\tau_{g}/2,s_{0,g}^{2}),\qquad Y_{g,i}(1)\sim N(b_{g}+\tau_{g}/2,s_{1,g}^{2})\\ &\tau=(\tau_{1},\tau_{2})^{\top}\sim N\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}\sigma_{1}^{2}&\rho\sigma_{1}\sigma_{2}\\ \rho\sigma_{1}\sigma_{2}&\sigma_{2}^{2}\end{pmatrix}\right),\qquad p(b_{g})\propto 1\end{split}$ (28)
Throughout, we set $s_{0,g}=s_{1,g}=1$, and $\sigma_{g}=0.1$ for all $g$, and investigate the change in the resulting optimal sample selection as we vary $\alpha_{1}$. As in Section 4.2, we determine the sample size using (22) based on a power calculation for the hypothesis test $H_{0}:\tau=0$ versus $H_{1}:\tau=0.1$, resulting in a total sample size of $N=1713$.
Figure 2: Bayes-optimal sample selection under Gaussian prior with $\rho=0$ and $\rho=0.9$, and under the least favorable prior (minimax-regret sample selection), along with the baseline proportional selection. Left: Proportion of group 1 in sample vs. in population. Right: Bayesian expected regret of the selections under Gaussian prior with $\rho=0$ and $\rho=0.9$, and under the least favorable prior.
Figure 2 displays (1) the Bayes-optimal sample selection with $\rho=0$, (2) the Bayes-optimal sample selection with $\rho=0.9$, (3) the minimax-regret sample selection, and their corresponding performances under three different priors. Given the specific prior and budget constraints used here, the Bayes-optimal selection under Gaussian prior with $\rho=0$ is close to the minimax-regret selection that oversamples the minority group, while the Bayes-optimal selection under Gaussian prior with $\rho=0.9$ is close to the proportional selection. [^7] As one would expect, using $\rho\in(0,0.9)$ here leads to selections in between proportional selection and minimax selection. When evaluated under the least favorable prior and the Gaussian prior with $\rho=0$, the minimax-regret selection and the Bayes-optimal selection with $\rho=0$ perform better than the other two selections, especially when the minority group has a small weight. Under the Gaussian prior with $\rho=0.9$, all selections show similar performances. As anticipated, the expected regret evaluated under the prior associated with the minimax-regret selection is consistently the highest, since it is the least favorable one.
These findings suggest that, in a simple two-group setting where the sample size is chosen such that the experiment has power for typical effect sizes under the prior, our minimax rules form a reasonable starting point for a discussion on sample selection: With uncorrelated priors ($\rho=0$) our recommendations closely match the Bayes-optimal ones, while with highly correlated priors ($\rho=0.9$), all methods do comparably well (because it’s possible to share information across groups). We do note, however, that it is also possible to construct examples where the Bayes-optimal rule diverges from our recommendations; for example, in the appendix we show a highly under-powered example in which the Bayes-optimal rule chooses to over-sample the majority group. Getting a deeper understanding of instance-level properties of the minimax-regret rules—and of when this approach is aligned with Bayesian sample selection with an informative prior—presents an interesting direction for further work.
@@ 6 Discussion
While utilitarian social welfare is a natural and common choice, it does favor groups with large proportion weights, which may sometimes be regarded as unfair. An alternative is to consider the egalitarian rule that suggests making decisions that maximize the minimum utility of all individuals in society under the least undesirable condition. Our findings demonstrate that this indeed results in equal sample allocation across all groups when the signal-to-noise ratio is the same across all groups. In scenarios where signal-to-noise ratios vary among groups, the resulting allocation suggests sampling proportionally to the variance of the treatment-control difference, ensuring an equal level of confidence in estimating the treatment effect across all groups.
@@
Finally, in this paper, we considered welfare-optimizing sample selection through the lens of minimax-regret decision rules. This approach is attractive in terms of its robustness and generality. However, in settings where prior evidence about the effectiveness of a treatment is available, it may be desirable to instead use Bayesian methods that incorporate this prior knowledge. Another interesting direction would be to investigate optimal sample selection under non-linear objectives. Models for cooperative bargaining can be used to motivate targeting expected log-utility (Nash, 1950; Kaneko and Nakamura, 1979); and Baek and Farias (2024) recently studied implications of targeting such objectives in bandit experiments. It may also be of interest to consider a richer welfare model that takes into account potentially different risks and benefits for study participants (who may risk direct harms from trying out a potentially dangerous treatment) and the general population (whose risks and benefits are mediated by the overall study findings).
@@ Acknowledgment
Appendix A A two-group normal-normal example
In this section, we derive the posterior and decision distributions for a normal-normal conjugate prior with $G=2$. Specifically, we consider the case where $p(\tau)\sim N(0,\Sigma)$ and treatments are assigned in a 1:1 completely stratified design with at most $N$ total units. We assume that $Y_{g,i}(0)$ and $Y_{g,i}(1)$ are generated i.i.d. from Gaussian distributions $N(b_{g}-\tau_{g}/2,s_{g}^{2})$ and $N(b_{g}+\tau_{g}/2,s_{g}^{2})$, respectively.
We start by deriving a closed-form expression of the distribution of $\hat{\tau}^{\text{PM}}$.
[…]
$\propto p(\tau)\int_{b}\prod_{g}\exp\left\{-\sum_{i:W_{g,i}=0}\frac{\left(Y_{g,i}-b_{g}+\tau_{g}/2\right)^{2}}{2s_{g}^{2}}\right\}\exp\left\{-\sum_{i:W_{g,i}=1}\frac{\left(Y_{g,i}-b_{g}-\tau_{g}/2\right)^{2}}{2s_{g}^{2}}\right\}db$
$\propto p(\tau)\int_{b}\prod_{g}\exp\left\{-\sum_{i:W_{g,i}=0}\frac{b_{g}^{2}-2b_{g}Y_{g,i}-b_{g}\tau_{g}+\tau_{g}Y_{g,i}+\tau_{g}^{2}/4}{2s_{g}^{2}}\right\}$
$\qquad\qquad\exp\left\{-\sum_{i:W_{g,i}=1}\frac{b_{g}^{2}-2b_{g}Y_{g,i}+b_{g}\tau_{g}-\tau_{g}Y_{g,i}+\tau_{g}^{2}/4}{2s_{g}^{2}}\right\}$
$\propto p(\tau)\prod_{g}\exp\left\{\frac{\tau_{g}\left(\sum_{i:W_{g,i}=1}Y_{g,i}-\sum_{i:W_{g,i}=0}Y_{g,i}\right)-n_{g}\tau_{g}^{2}/4}{2s_{g}^{2}}\right\}$
$\qquad\qquad\int_{b}\prod_{g}\exp\left\{-\frac{n_{g}}{2s_{g}^{2}}\cdot b_{g}^{2}+\frac{\sum_{i:W_{g,i}=0}Y_{g,i}+\sum_{i:W_{g,i}=1}Y_{g,i}}{s_{g}^{2}}\cdot b_{g}\right\}db$
$\propto p(\tau)\prod_{g}\exp\left\{\frac{\tau_{g}\left(\sum_{i:W_{g,i}=1}Y_{g,i}-\sum_{i:W_{g,i}=0}Y_{g,i}\right)-n_{g}\tau_{g}^{2}/4}{2s_{g}^{2}}\right\}$
$\propto p(\tau)\prod_{g}\exp\left\{-\frac{n_{g}}{8s_{g}^{2}}\left(\tau_{g}^{2}-2\cdot\frac{\sum_{i:W_{g,i}=1}Y_{g,i}-\sum_{i:W_{g,i}=0}Y_{g,i}}{n_{g}/2}\cdot\tau_{g}\right)\right\}.$ (S1)
Note that the production term in (S1) above is the kernel of a multivariate Gaussian distribution with mean
$\mu_{D}=\left(\frac{\sum_{i:W_{1,i}=1}Y_{1,i}-\sum_{i:W_{1,i}=0}Y_{1,i}}{n_{1}/2},\frac{\sum_{i:W_{2,i}=1}Y_{2,i}-\sum_{i:W_{2,i}=0}Y_{2,i}}{n_{2}/2}\right)^{\top},$ (S2)
and variance
$\Sigma_{D}=\operatorname{diag}\left(\frac{4s_{1}^{2}}{n_{1}},\frac{4s_{2}^{2}}{n_{2}}\right).$ (S3)
Thus, by standard result from Bayesian inference with conjugate priors [Gelman et al., 2013],
$\hat{\tau}^{\text{PM}}=\left(\Sigma^{-1}+\Sigma_{D}^{-1}\right)^{-1}\Sigma_{D}^{-1}\mu_{D}.$ (S4)
Since
$\begin{split}&\mathbb{E}\left[\frac{2}{n_{g}}\sum_{i=1}^{n_{g}}W_{g,i}Y_{g,i}-\frac{2}{n_{g}}\sum_{i=1}^{n_{g}}(1-W_{g,i})Y_{g,i}\,\big{|}\,\tau_{g}\right]\\ &\qquad\qquad=\mathbb{E}\left[Y_{g,i}\,\big{|}\,W_{g,i}=1,\tau_{g}\right]-\mathbb{E}\left[Y_{g,i}\,\big{|}\,W_{g,i}=0,\tau_{g}\right]\\ &\qquad\qquad=\tau_{g},\end{split}$ (S5)
and
$\begin{split}&\operatorname{Var}\left[\frac{2}{n_{g}}\sum_{i=1}^{n_{g}}W_{g,i}Y_{g,i}-\frac{2}{n_{g}}\sum_{i=1}^{n_{g}}(1-W_{g,i})Y_{g,i}\,\big{|}\,\tau_{g}\right]\\ &\qquad\qquad=\operatorname{Var}\left[\mathbb{E}\left[\frac{2}{n_{g}}\sum_{i=1}^{n_{g}}W_{g,i}Y_{g,i}-\frac{2}{n_{g}}\sum_{i=1}^{n_{g}}(1-W_{g,i})Y_{g,i}\,\big{|}\,\tau_{g},W_{1:n_{g}}\right]\,\big{|}\,\tau_{g}\right]+\\ &\qquad\qquad\qquad\qquad\mathbb{E}\left[\operatorname{Var}\left[\frac{2}{n_{g}}\sum_{i=1}^{n_{g}}W_{g,i}Y_{g,i}-\frac{2}{n_{g}}\sum_{i=1}^{n_{g}}(1-W_{g,i})Y_{g,i}\,\big{|}\,\tau_{g},W_{1:n_{g}}\right]\,\big{|}\,\tau_{g}\right]\\ &\qquad\qquad=\operatorname{Var}\left[\frac{2}{n_{g}}\sum_{i=1}^{n_{g}}W_{g,i}(b_{g}+\tau_{g})-\frac{2}{n_{g}}\sum_{i=1}^{n_{g}}(1-W_{g,i})b_{g}\,\big{|}\,\tau_{g}\right]+\\ &\qquad\qquad\qquad\qquad\mathbb{E}\left[\frac{4}{n_{g}^{2}}\sum_{i=1}^{n_{g}}W_{g,i}s_{g}^{2}+\frac{4}{n_{g}^{2}}\sum_{i=1}^{n_{g}}W_{g,i}s_{g}^{2}\,\big{|}\,\tau_{g}\right]\\ &\qquad\qquad=\operatorname{Var}\left[\frac{2}{n_{g}}\sum_{i=1}^{n_{g}}(2W_{g,i}-1)b_{g}+\frac{2}{n_{g}}\sum_{i=1}^{n_{g}}W_{g,i}\tau_{g}\,\big{|}\,\tau_{g}\right]+\\ &\qquad\qquad\qquad\qquad\mathbb{E}\left[\frac{4}{n_{g}^{2}}\sum_{i=1}^{n_{g}}W_{g,i}s_{g}^{2}+\frac{4}{n_{g}^{2}}\sum_{i=1}^{n_{g}}W_{g,i}s_{g}^{2}\,\big{|}\,\tau_{g}\right]\\ &\qquad\qquad=\frac{4s_{g}^{2}}{n_{g}},\end{split}$ (S6)
where the last equality holds due to that $\sum_{i=1}^{n_{g}}W_{g,i}=n_{g}/2$ in a stratified completely randomized experiment with units assigned treatment or control in a 1:1 ratio,
$\hat{\tau}^{\text{PM}}\,\big{|}\,\tau\sim N\left(\left(\Sigma^{-1}+\Sigma_{D}^{-1}\right)^{-1}\Sigma_{D}^{-1}\tau,\left(\Sigma^{-1}+\Sigma_{D}^{-1}\right)^{-1}\Sigma_{D}^{-1}\left(\Sigma^{-1}+\Sigma_{D}^{-1}\right)^{-1}\right).$ (S7)
Thus, it is possible to calculate directly $\mathbb{E}\left[I(\hat{\tau}^{\text{PM}}_{g}(D;n,\tau)>0)\,\big{|}\,\tau_{g}\right]=\mathbb{P}\left[\hat{\tau}^{\text{PM}}_{g}(D;n,\tau)>0\,\big{|}\,\tau_{g}\right]$ using the cdf of a standard Gaussian distribution.
To calculate the expected regret, we note that
$\begin{split}H^{p}(\mathbf{n})&=\mathbb{E}_{\tau\sim p}\left[\mathbb{E}_{D}\left[R(\mathbf{n},\delta^{p})\,\big{|}\,\tau\right]\right]\\ &=\sum_{g}\alpha_{g}\mathbb{E}_{\tau\sim p}\left[\tau_{g}I(\tau_{g}>0)-\tau_{g}\mathbb{P}_{D}\left[\hat{\tau}^{\text{PM}}_{g}>0\,\big{|}\,\tau_{g}\right]\right].\end{split}$ (S8)
From (S7), it is possible to obtain $\mathbb{P}_{D}\left[\hat{\tau}^{\text{PM}}_{g}>0\,\big{|}\,\tau_{g}\right]$ using the cdf of a standard Gaussian distributional function.
Appendix B Additional Simulation Example for Bayes-Optimal Sample Selection
In this section, we provide additional experiments and extend the simulation in Section 5 to a highly under-powered setting. We use the same data-generating distribution detailed in Section 5, but with a much smaller total sample size of $N=100$. Figure S1 displays (1) the Bayes-optimal sample selection with $\rho=0$, (2) the Bayes-optimal sample selection with $\rho=0.9$, (3) the minimax-regret sample selection, and their corresponding performances under three different priors.
In contrast to the minimax-regret selection, the Bayes-optimal selections here oversamples the majority group. This is because—given the specific prior and the small budget—attempting to learn from both groups may lead to suboptimal decisions for the entire population, whereas allocating more samples to the majority group ensures that at least a significant portion of the population is treated appropriately. Moreover, we observe increased oversampling of the majority group with greater data sharing, in which case allocating sufficient samples to one group can benefit decision-making for the entire population. We emphasize, however, that if the chosen prior is misspecified, the sample selections induced by the prior could result in large or even extreme regret in an adversarial environment. For example, in the right panel of Figure S1, we see that the expected regret of the $\rho=0.9$ Bayes-optimal rule explodes when evaluated against the least-favorable prior that underlies our minimax-regret rule.
Figure S1: Bayes-optimal sample selection under Gaussian prior with $\rho=0$ and $\rho=0.9$, and under the least favorable prior (minimax-regret sample selection), along with the baseline proportional selection. Left: Proportion of group 1 in sample vs. in population. Right: Bayesian expected regret of the selections under Gaussian prior with $\rho=0$ and $\rho=0.9$, and under the least favorable prior.
Appendix C Proof of Theorems
C.1 Proof of Theorem 15